# Heterogeneous Intertemporal Treatment Effects via Dynamic Panel Data Models

### v1

@@ 5 Conclusion
In this note we considered the causal interpretation of dynamic panel data estimators in models of potential outcomes that can depend on past treatments. We derived sufficient conditions for instrumented first-difference estimators to be causally interpretable and proposed a new adjusted inverse probability of treatment estimator that uses the first-differenced outcome to recover a history-dependent average treatment effect.
An underlying assumption for (most of) these results was sequential exchangeability. In addition to investigating the plausibility of this assumption in a structural model of dynamic choice, we showed that this assumption is not necessary per se for dynamic panel data estimators to recover causal estimands in our setting. A natural avenue for future work is to consider (partial) identification when sequential exchangeability is relaxed or holds only conditional on a set of unobservables.
Appendix A: Proofs
Proof of Lemma 1. Let $\mathcal{Y}$ denote the (discrete) support of $Y_{0}$. Let $Z_{i}\equiv\{\mathbb{I}[(D_{i1},Y_{i0})=(d_{1},y_{0})]\}_{d_{1}\in\{0,1\},y_{0}\in\mathcal{Y}}$ be a row-vector of $2\times\#\mathcal{Y}$ (dummy) instrumental variables. The 2SLS estimator for the coefficient of $\Delta D_{i2}$ in (8) is constructed as follows. In the (fully saturated) first stage, obtain the fitted values of $\Delta Y_{i1}$ (and $\Delta D_{i2}$) from their regression on $Z_{i}$. These fitted values are $Z_{i}\hat{A}$ (and $Z_{i}\hat{B}$, respectively), where
$\hat{A}\equiv\left(\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}Z_{i}\right)^{-1}\left[\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}\left(\Delta Y_{i1}\right)\right]\text{ and }\hat{B}\equiv\left(\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}Z_{i}\right)^{-1}\left[\frac{1}{n}\sum\nolimits_{i}Z_{i}^{\prime}\left(\Delta D_{i2}\right)\right]\text{.}$
By the property of ordinary least squares estimators,
[…]
Formally, $A$ (and $B$) are $(2\times\#\mathcal{Y)}$-by-$1$ column vectors with each component being the conditional mean of $\Delta Y_{i1}$ (and $\Delta D_{i2}$) given $(D_{i1},Y_{i0})=(d_{1},y_{0})$ for some $d_{1}\in\{0,1\}$, $y_{0}\in\mathcal{Y}$. This means $Z_{i}A=E[\Delta Y_{i1}|D_{i1},Y_{i0}]$ and $Z_{i}B=E[\Delta D_{i2}|D_{i1},Y_{i0}]$ by construction.
In the second stage, regress $\Delta Y_{i2}$ on an intercept and the first-stage fitted values $Z_{i}\hat{A}$, $Z_{i}\hat{B}$. Our goal is to investigate the causal interpretation of the slope estimate for $Z_{i}\hat{B}$ in this second-stage regression. By the Frisch-Waugh-Lovell Theorem, this slope estimate is numerically equivalent to
$\frac{\frac{1}{n}\sum_{i}\hat{W}_{i}\Delta Y_{i2}}{\frac{1}{n}\sum_{i}\hat{W}_{i}^{2}}\text{,}$ (36)
where $\hat{W}_{i}$ is the residual from regressing $Z_{i}\hat{B}$ on $[1,Z_{i}\hat{A}]\equiv\hat{M}_{i}$. That is,
$\hat{W}_{i}$ $=$ $Z_{i}\hat{B}-\hat{M}_{i}\underset{\equiv\hat{C}}{\underbrace{\left(\frac{1}{n}\sum\nolimits_{i}\hat{M}_{i}^{\prime}\hat{M}_{i}\right)^{-1}\left(\frac{1}{n}\sum\nolimits_{i}\hat{M}_{i}^{\prime}Z_{i}\right)}}\hat{B}$ (37)
$=$ $\left(Z_{i}-\hat{M}_{i}\hat{C}\right)\hat{B}\text{,}$
where
[…]
We now turn to the denominator of the 2SLS estimator of slope coefficient of $\Delta D_{2}$ in (36):
$\frac{1}{n}\sum\nolimits_{i}\hat{W}_{i}^{2}=$ $\hat{B}^{\prime}\left[\frac{1}{n}\sum\nolimits_{i}(Z_{i}-\hat{M}_{i}\hat{C})^{\prime}(Z_{i}-\hat{M}_{i}\hat{C})\right]\hat{B}$
[…]
$=$ $E\left[(Z_{i}B-M_{i}CB)^{\prime}(Z_{i}B-M_{i}CB)\right],$
[…]
Next, note the numerator is
$\frac{1}{n}\sum\nolimits_{i}\hat{W}_{i}\left(\Delta Y_{i2}\right)=$ $\hat{B}^{\prime}\left[\frac{1}{n}\sum\nolimits_{i}(Z_{i}-\hat{M}_{i}\hat{C})^{\prime}(\Delta Y_{i2})\right]$
[…]
$=$ $E\left[(Z_{i}B-M_{i}CB)^{\prime}(\Delta Y_{i2})\right]=E[w(Y_{i0},D_{i1})\left(\Delta Y_{i2}\right)]\text{,}$
[…]
Proof of Proposition 1. Throughout the following proof, we fix $s\geq 1$. For $t=s$, the equivalence is immediate:
@@
Finally, recall we have already shown the equivalence (between the two sets of conditions on the levels and the trends of potential outcomes) for $t=s$ at the beginning of this proof. Therefore, the claim of the reverse direction of implication for $t>s$ holds.
@@
Proof of Proposition 2. By Lemma 1, the estimator for the coefficient $\beta$ converges in probability to (13), recalled here for convenience:
$E\left[\frac{w(Y_{0},D_{1})D_{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right]+E\left[\frac{w(Y_{0},D_{1})}{E[w(Y_{0},D_{1})^{2}]}\times\delta_{2}(D_{1})\right]$
where $w(Y_{0},D_{1})$ is the projection error in the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|Y_{0},D_{1})$.
The first step of the proof is to derive a closed-form expression for the projection errors $w(Y_{0},D_{1})$. We have:
@@
where the first equality is due to the decomposition in (11), the second equality rearranges, the third equality follows from mean sequential exchangeability (Assumption 2b), specifically (40), and the fourth equality follows from limited heterogeneity (Assumption 3), specifically (14) and (16). Furthermore, by (16), this expression is strictly monotone in the realization $D_{1}$. Therefore, the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|D_{1},Y_{0})$ is identical to its mean conditional on $D_{1}$. This implies:
@@
The second step of the proof is to eliminate the second bias term in (13) by establishing that its numerator:
@@
equals zero. Applying the law of iterated expectations (LIE) over partitions $(Y_{0},D_{1})$, this numerator equals:
$E[w(Y_{0},D_{1})E[\delta_{2}(D_{1})|Y_{0},D_{1}]]$ (44)
Focusing on the inner expectation for treatment realization $d_{1}$,
@@
The third step of the proof is to establish the convex weights over heterogeneous treatment effects in the first term of (13), with numerator $E[w(Y_{0},D_{1})D_{2}\tau_{2}(D_{1})]$. Applying the LIE over partitions $(Y_{0},Y_{1},D_{1},D_{2})$ and invoking mean sequential exchangeability (Assumption 2b), specifically (42), yields:
@@
where the first equality follows by definition of the observed variables at realized treatment, the second by the full sequential exchangeability (Assumption 2a) for period 1 in (38), specifically its implication that $D_{1}\perp Y_{2}(\cdot)|Y_{0},Y_{1}(d_{1})$ for any $d_{1}$, and the third by limited heterogeneity (Assumption 3), specifically (17). Similarly to the previous step, we also have:
@@
$=E[\tau_{2}(d_{1})]$
@@
Proceeding with expression (47) for the first numerator:
@@
where the first equality follows from (49), the second by the LIE over partition $(D_{1},Y_{0})$, the third by the LIE and definition of $w(Y_{0},D_{1})$ (which imply $E[w(Y_{0},D_{1})g(D_{1})]=0$ for any function $g(\cdot)$ with a sole argument $D_{1}$), the fourth by the previously derived projection errors (43), the fifth again by (49), the sixth by rearrangement, and the seventh again by the LIE.
@@
$E[\widetilde{\Delta Y_{2}}]=E\left[\frac{D_{2}\tau_{2}(D_{1})+\delta_{2}(D_{1})-E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]}{E[D_{2}|Y^{1},D_{1}]}\right]$ (50)
@@
where the first equality follows by definition of $\Delta Y_{2}$, the second by plugging in observed variables and rearrangement, the third by mean sequential exchangeability (Assumption 2), specifically (42), and the fourth by rearrangement and the definition (9). Iterating expectations over the previous equality,
@@
where the first equality follows by the law of iterated expectations (LIE) over $(Y^{1},D_{1})$, the second by (51), the third again by the LIE over $(Y^{1},D_{1},D_{2})$, the fourth by mean sequential exchangeability (Assumption 2), specifically (42), the fifth by rearrangement and cancellation of terms $E[D_{2}|Y^{1},D_{1}]$, and the last again by the LIE. (Note mean sequential exchangeability is invoked only for $D_{2}$ in (42).)
Proof of Proposition 4. For all $s\geq 1$,
$E(\varepsilon_{is}\alpha_{i})$ $=$ $E[\alpha_{i}E\left(\varepsilon_{is}|Y_{i0},\alpha_{i}\right)]$
$=$ $E\left\{\alpha_{i}\left[\sum\nolimits_{d^{s}}E\left(\varepsilon_{is}|D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right)\times\Pr\{D_{i}^{s}=d^{s}|Y_{i0},\alpha_{i}\}\right]\right\}$
$=$ $E\left\{\alpha_{i}\left[\sum\nolimits_{d^{s}}E\left(\varepsilon_{is}^{\ast}(d_{s})|D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right)\times\Pr\{D_{i}^{s}=d^{s}|Y_{i0},\alpha_{i}\}\right]\right\}=0\text{,}$
where the first and second equalities are due to the law of iterated expectation, the third due to the definition $\varepsilon_{it}=\varepsilon_{it}^{\ast}(D_{it})$, and the last due to the sequential exogeneity of $\varepsilon_{it}^{\ast}(\cdot)$ in (25). Hence $E(\Delta\varepsilon_{it}\alpha_{i})=0$ for all $t\geq 2$. Likewise, swapping the roles of $\alpha_{i}$ and $Y_{i0}$ and applying the same argument above shows that $E(\Delta\varepsilon_{it}Y_{i0})=0$.
Furthermore, note that for $s\leq t$, the law of iterated expectation implies
$E(D_{is}\varepsilon_{it})=E[D_{is}E(\varepsilon_{it}|D_{i}^{t},Y_{i0},\alpha_{i})]\text{,}$
where
$E\left[\varepsilon_{it}|D_{i}^{t}=d^{t}\equiv(d^{t-1},d_{t}),Y_{i0},\alpha_{i}\right]$
$=$ $E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=(d^{t-1},d_{t}),Y_{i0},\alpha_{i}\right]=0$
for all $d^{t}$, where the second equality holds due to (25). Thus (25) implies (28).
Next, without loss of generality, consider $s<t$ so that
$E(\varepsilon_{is}\varepsilon_{it})$ $=$ $E[E(\varepsilon_{is}\varepsilon_{it}|\varepsilon_{is},D_{i}^{s},Y_{i0},\alpha_{i})]$
$=$ $E\left\{\varepsilon_{is}^{\ast}(D_{is})E[\varepsilon_{it}|\varepsilon_{is}^{\ast}(D_{is}),D_{i}^{s},Y_{i0},\alpha_{i}]\right\}\text{.}$
By the law of iterated expectations,
$E\left[\varepsilon_{it}|\varepsilon_{is}^{\ast}(D_{is}),D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\right]$
$=$ $\sum_{d^{t}\in\omega_{t}(d^{s})}E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=d^{t},\varepsilon_{is}^{\ast}(d_{s}),Y_{i0},\alpha_{i}\right]\Pr\{D_{i}^{t}=d^{t}|\varepsilon_{is}^{\ast}(d_{s}),D_{i}^{s}=d^{s},Y_{i0},\alpha_{i}\}$
$=$ $0\text{,}$
where $\omega_{t}(d^{s})$ denotes the support of $D_{i}^{t}$ conditional on $D_{i}^{s}=d^{s}$, and the last equality holds because $E\left[\varepsilon_{it}^{\ast}(d_{t})|D_{i}^{t}=d^{t},\varepsilon_{is}^{\ast}(d_{s}),Y_{i0},\alpha_{i}\right]$ is zero under (25).
In what follows, it will be useful to collect the following relationships between potential outcomes and composite errors $\nu_{it}^{*}(d_{t})\equiv\alpha_{i}^{*}+\varepsilon_{it}^{*}(d_{t})$ in the linear dynamic model of potential outcomes (21).
@@
For any $d^{t}\in\{0,1\}^{t}$ and $t\geq s\geq 1$,
$\nu_{it}^{*}(d_{t})$ $=Y_{it}(d^{t})-\gamma^{*}Y_{it-1}(d^{t-1})-\beta^{*}d_{t}-\theta_{t}^{*}$ (52)
$Y_{it}(d^{t})$ $=(\gamma^{*})^{t-s+1}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]$ (53)
@@
$Y_{it-1}(d^{t-1})=(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]$
@@
$Y_{it}(d^{t})=\gamma^{*}Y_{it-1}(d^{t-1})+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d_{t})$
@@
$Y_{it}(d^{t})=$ $\gamma^{*}\left[(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d_{r})]\right]$
$+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d_{t})$
@@
$\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\alpha_{i}^{*}+E[\varepsilon_{ir}^{*}(d_{r})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]]$
@@
$E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]$
@@
$E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\nu_{i}^{*s-1}(d^{s-1})]$
@@
$E[\nu_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon_{i}^{*s-1}(d^{s-1})]$
@@
$E[E[\varepsilon_{it}^{*}(d_{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{t},\varepsilon^{*t-1}(\cdot)]|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon^{*s-1}(d^{s-1})]=0.$
Proof of Proposition 6.
Fix a sequence of treatment $d^{T}$. For simplicity of notation, expand a vector-valued representation of choice rule (28) as:
$D^{t}=\phi^{t}(D^{t-1},Y^{t-1}(D^{t-1}),\zeta^{t}(D^{t-1}))$
so that we can express the event:
$\{D^{t}=d^{t}\}=\{\phi^{t}(d^{t-1},Y^{t-1}(d^{t-1}),\zeta^{t}(d^{t-1}))=d^{t}\}$ (54)
Condition (29) implies:
$Y^{t}(d^{t})\perp\zeta^{t}(d^{t-1})\mid Y_{0}$
because $Y^{t}(d^{t})$ is a function of $(Y_{0},\alpha(\cdot),\varepsilon^{t}(d^{t}))$ by definition (27). By rules of conditional independence,
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}(d^{t-1})\mid Y^{t-1}(d^{t-1})$
and furthermore, for $t\geq 2$,
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}(d^{t-1})\mid Y^{t-1}(d^{t-1}),\,\phi^{t-1}(d^{t-2},Y^{t-2}(d^{t-2}),\zeta^{t-1}(d^{t-2}))=d^{t-1}$
because, having conditioned on $Y^{t-1}(d^{t-1})$, the second conditioning statement is a function of $\zeta^{t-1}(d^{t-2})$. By (54) and because $\zeta^{t}(D^{t-1})=\zeta^{t}$ and $Y^{t-1}(D^{t-1})=Y^{t-1}$, this reduces to:
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\zeta^{t}\mid Y^{t-1},\,D^{t-1}=d^{t-1}$
Conditioning on history $(D^{t-1},Y^{t-1})$, this implies that, for a deterministic function $\phi_{t}(\cdot)$:
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp\phi_{t}(D^{t-1},Y^{t-1},\zeta^{t})\mid Y^{t-1},\,D^{t-1}=d^{t-1}$
which by definition (28) yields the desired result:
$(Y_{t}(d^{t}),\dots,Y_{T}(d^{T}))\perp D_{t}|Y^{t-1},D^{t-1}=d^{t-1}$
In the case where $t=1$, the conditioning event collapses to $Y_{0}$ and the argument proceeds analogously.
Proof of Proposition 7. Assumption 2b holds under (33) by the same argument as in the proof of Proposition 6. We adapt it here for the weaker case of mean sequential exchangeability in simpler notation. With potential outcomes and treatment rules in (30) and (31), we have:

### v2

@@ 5 Conclusion
We studied the causal interpretation of dynamic panel data estimators in models where potential outcomes depend on past treatments. We derived sufficient conditions for instrumented first-difference estimators to be causally interpretable. A motivation and underlying assumption for one set of these results was sequential exchangeability. Besides investigating the plausibility of this assumption in both experimental and observational settings (with a structural model of dynamic treatment choices), we showed how to conduct causal inference using dynamic panel data estimators when this assumption fails. A natural avenue for future work is to consider (partial) identification when sequential exchangeability is relaxed or holds only conditional on a set of unobservables.
Appendix A Proofs
Proof of Proposition 1. We fix $s\geq 1$ throughout the following proof. For $t=s$, the equivalence is immediate:
@@ Appendix A Proofs
Finally, recall we already showed the equivalence (between the two sets of conditions on the levels and the trends of potential outcomes) for $t=s$ at the start of this proof. Hence by induction the claim of the reverse direction of implication for $t>s$ holds.
[…]
$E[Z^{\prime}\Delta X]^{-1}E[Z^{\prime}\Delta Y_{2}]$
Note that $Z$ contains conditional expectations of each element of $\Delta X$. By the law of iterated expectations, this estimand is equal to an OLS estimand:
$E[Z^{\prime}Z]^{-1}E[Z^{\prime}\Delta Y_{2}].$
By the formula for partitioned regressions, a.k.a. a population version of the Frisch-Waugh-Lovell Theorem, the estimand for the slope coefficient of the first component in $Z$ above is equal to
$\frac{E[w(Y_{0},D_{1})\Delta Y_{2}]}{E[w(Y_{0},D_{1})^{2}]},$
where $w(Y_{0},D_{1})$, defined in (9), is the projection error from the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ onto $E(\Delta Y_{1}|Y_{0},D_{1})$ and a constant intercept. Plugging in the decomposition (8) of $\Delta Y_{2}$ yields the desired expression in (10).
@@
Proof of Proposition 2. By Lemma 1, the (first-differenced) IV estimand $\widetilde{\beta}$ is:
$E\left[\frac{w(Y_{0},D_{1})D_{2}}{E[w(Y_{0},D_{1})^{2}]}\times\tau_{2}(D_{1})\right]+E\left[\frac{w(Y_{0},D_{1})}{E[w(Y_{0},D_{1})^{2}]}\times\delta_{2}(D_{1})\right],$ (41)
where $w(Y_{0},D_{1})$ is the error in the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|Y_{0},D_{1})$.
First, we derive a closed-form expression for the projection errors $w(Y_{0},D_{1})$. We have:
@@
where the first equality is due to the decomposition in (7), the second equality rearranges, the third equality follows from mean sequential exchangeability (Assumption 2b), specifically (38), and the fourth equality follows from limited heterogeneity (Assumption 3), specifically (11) and (13). Furthermore, by (13), this right-hand side above is strictly monotone in $D_{1}$. Therefore, the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|D_{1},Y_{0})$ is identical to its mean conditional on $D_{1}$. This implies:
@@
Next, we eliminate the second (bias) term in (41) by showing its numerator:
@@
is zero. Applying the law of iterated expectations (LIE) over partitions $(Y_{0},D_{1})$, we write this numerator as:
$E[w(Y_{0},D_{1})E[\delta_{2}(D_{1})|Y_{0},D_{1}]].$ (43)
The inner expectation conditional on a realized treatment $D_{1}=d_{1}$ equals:
@@
The last step of the proof is to establish the convex weights over heterogeneous treatment effects in the first term of (41), with numerator $E[w(Y_{0},D_{1})D_{2}\tau_{2}(D_{1})]$. Applying the LIE over partitions $(Y_{0},Y_{1},D_{1},D_{2})$ and invoking mean sequential exchangeability (Assumption 2b), specifically (40), yields:
@@
where the first equality follows by definition of the observed variables at realized treatment, the second by the full sequential exchangeability (Assumption 2a) for period 1 in (36), specifically its implication that $D_{1}\perp Y_{2}(\cdot)|Y_{0},Y_{1}(d_{1})$ for any $d_{1}$, and the third by limited heterogeneity (Assumption 3), specifically (14). Similarly to the previous step, we have:
@@
$=E[\tau_{2}(d_{1})],$
where the first equality follows from the LIE, the second from (47), and the third from taking an expectation over a constant. Combining with (47), it follows that:
@@
Thus, the right-hand side of (46) is:
@@
where the first equality follows from (48), the second by the LIE over partition $(D_{1},Y_{0})$, the third by the LIE and definition of $w(Y_{0},D_{1})$ (which implies $E[w(Y_{0},D_{1})g(D_{1})]=0$ for any function $g(\cdot)$ with a sole argument $D_{1}$), the fourth by the previously derived projection errors (42), the fifth again by (48), the sixth by rearrangement, and the seventh again by the LIE.
@@
$E[\widetilde{\Delta Y_{2}}]=E\left[\frac{D_{2}\tau_{2}(D_{1})+\delta_{2}(D_{1})-E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]}{E[D_{2}|Y^{1},D_{1}]}\right].$ (49)
@@
where the first equality follows by definition of $\Delta Y_{2}$, the second by plugging in observed variables and rearrangement, the third by mean sequential exchangeability (Assumption 2b), specifically (40), and the fourth by rearrangement and the definition of $\delta_{2}$ in (6). Iterating expectations over the previous equality,
@@
where the first equality follows by the law of iterated expectations (LIE) over $(Y^{1},D_{1})$, the second by (50), the third again by the LIE over $(Y^{1},D_{1},D_{2})$, the fourth by mean sequential exchangeability (Assumption 2b), specifically (40), the fifth by rearrangement and cancellation of terms $E[D_{2}|Y^{1},D_{1}]$, and the last again by the LIE. (Note mean sequential exchangeability is invoked only for $D_{2}$ in (40).)
In what follows, it will be useful to collect the following relationships between potential outcomes and composite errors $\nu_{it}^{*}(d^{t})\equiv\alpha_{i}^{*}+\varepsilon_{it}^{*}(d^{t})$ in the linear dynamic model of potential outcomes (17).
@@
Suppose the potential outcomes follow an AR(1) model in (17). For any $d^{t}\in\{0,1\}^{t}$ and $t\geq s\geq 1$,
$\nu_{it}^{*}(d^{t})$ $=Y_{it}(d^{t})-\gamma^{*}Y_{it-1}(d^{t-1})-\beta^{*}d_{t}-\theta_{t}^{*},$ (51)
$Y_{it}(d^{t})$ $=(\gamma^{*})^{t-s+1}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d^{r})].$ (52)
@@
$Y_{it-1}(d^{t-1})=(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d^{r})]$
@@
$Y_{it}(d^{t})=\gamma^{*}Y_{it-1}(d^{t-1})+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d^{t})$
@@
$Y_{it}(d^{t})=$ $\gamma^{*}\left[(\gamma^{*})^{t-s}Y_{is-1}(d^{s-1})+\sum_{r=s}^{t-1}(\gamma^{*})^{t-1-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\nu_{ir}^{*}(d^{r})]\right]$
$+\beta^{*}d_{t}+\theta_{t}^{*}+\nu_{it}^{*}(d^{t})$
@@
$\sum_{r=s}^{t}(\gamma^{*})^{t-r}\times[\beta^{*}d_{r}+\theta_{r}^{*}+\alpha_{i}^{*}+E[\varepsilon_{ir}^{*}(d^{r})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]]$
@@
$E[\varepsilon_{it}^{*}(d^{t})|\alpha_{i}^{*},Y_{i}^{s-1}(d^{s-1}),D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime}]$
@@
$E[\varepsilon_{it}^{*}(d^{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\nu_{i}^{*s-1}(d^{s-1})]$
@@
$E[\varepsilon_{it}^{*}(d^{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon_{i}^{*s-1}(d^{s-1})]$
@@
$E[E[\varepsilon_{it}^{*}(d^{t})|\alpha_{i}^{*},Y_{i0},D_{i}^{t},\varepsilon^{*t-1}(\cdot)]|\alpha_{i}^{*},Y_{i0},D_{i}^{s-1}=d^{s-1},D_{is}=d_{s}^{\prime},\varepsilon^{*s-1}(d^{s-1})]=0.$
Proof of Proposition 5. Per Observation 1, when the DGP of potential outcomes satisfies Assumption 5, the implied sequence of observables $(D_{i}^{t},Y_{i}^{t})$ satisfies the reduced-form model of observed outcomes in (3) with $(\gamma,\beta,\theta_{t},\alpha_{i})=(\gamma^{*},\beta^{*},\theta_{t}^{*},\alpha_{i}^{*})$ and reduced-form errors $\varepsilon_{it}=\varepsilon_{it}^{*}(D_{i}^{t})$. First-differencing of (3) yields
$\Delta Y_{it}=\beta^{*}\Delta D_{it}+\gamma^{*}\Delta Y_{it-1}+\Delta\theta_{t}^{*}+\Delta\varepsilon_{it},$
where $E(\varepsilon_{it}|D_{i}^{t},Y_{i}^{t-1})=0$ because of the condition (18) in Assumption 5 and the fact that $Y_{i}^{t-1}$ is a function of $\{D_{i}^{t-1},Y_{i0},\varepsilon_{i}^{*t-1}(\cdot),\alpha_{i}\}$. Consequently, $E(\Delta\varepsilon_{it}|D_{i}^{t-1},Y_{i}^{t-2})=0$ for all $t\geq 2$. Using $E(\Delta D_{it}|D_{i}^{t-1},Y_{i}^{t-2})$ and $E(\Delta Y_{it-1}|D_{i}^{t-1},Y_{i}^{t-2})$ as instruments for $\Delta D_{it}$ and $\Delta Y_{it-1}$ in first-differenced instrumental-variable regression proves the result.
Proof of Proposition 6. Under Assumption 6, we use the first-differenced observed equation (23) to write the conditional mean of $\Delta Y_{it}$ given $W_{it-1},\mathcal{H}_{t-1}(y_{0})$ as:
$E[\Delta Y_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]=E[\tau^{*}_{it}D_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]+\gamma^{*}E[\Delta Y_{it-1}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]+\Delta\theta_{t}^{*},$ (53)
because $E(\Delta\varepsilon^{*}_{it}|H_{it-1})=0$ and $W_{it-1}$ is a function of $H_{it-1}$. Moreover,
$E[\tau^{*}_{it}D_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]$ $=$ $E[\tau^{*}_{it}(D_{i}^{t-1})|D_{it}=1,W_{it-1},\mathcal{H}_{t-1}(y_{0})]\times E[D_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]$
$=$ $\underset{ATFT_{t}(y_{0})}{\underbrace{E[\tau^{*}_{it}(D_{i}^{t-1})|D_{it}=1,\mathcal{H}_{t-1}(y_{0})]}}\times E[\Delta D_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})],$
where the first equality holds by the law of iterated expectation, and the second holds under Assumption 7 because $W_{it-1}$ is a function of $\{\varepsilon_{i}^{*t-1}(\cdot),\alpha^{*}_{i}\}$ and $D_{i}^{t-1}=0$ once conditional on $\mathcal{H}_{t-1}(y_{0})$. This implies the mean of $\Delta Y_{it}$ conditional on $W_{it-1}$ and the event $\mathcal{H}_{t-1}(y_{0})$ is linear in $E[Z_{it}^{\prime}(y_{0})|\mathcal{H}_{t-1}(y_{0})]$, with the coefficients being $(ATFT_{t}(y_{0}),\gamma^{*},\Delta\theta_{t}^{*})$. Therefore,
$(ATFT_{t}(y_{0}),\gamma^{*},\Delta\theta_{t}^{*})^{\prime}$ $=$ $E[Z_{it}^{\prime}Z_{it}|\mathcal{H}_{t-1}(y_{0})]^{-1}E[Z_{it}^{\prime}\Delta Y_{it}|\mathcal{H}_{t-1}(y_{0})]$
$=$ $E[Z_{it}^{\prime}X_{it}|\mathcal{H}_{t-1}(y_{0})]^{-1}E[Z_{it}^{\prime}\Delta Y_{it}|\mathcal{H}_{t-1}(y_{0})],$
where we suppress the dependence of instruments $Z_{it}(y_{0})$ on the initial condition $y_{0}$ for simplicity, and the second equality holds because $E[Z_{it}^{\prime}X_{it}|\mathcal{H}_{t-1}(y_{0})]=E[Z_{it}^{\prime}Z_{it}|\mathcal{H}_{t-1}(y_{0})]$ by the law of iterated expectation.
Proof of Proposition 8. With potential outcomes and treatment rules in (31) and (32), we have: